5.015 g sample of a compound that contained carbon, hydrogen, and oxygen was combusted in a carbon-hydrogen analyzer. The combustion produced 7.35 g of carbon dioxide and
2.99 g of water. Determine the empirical formula of the compound.
1
Expert's answer
2020-10-14T05:45:53-0400

Let the compound be CxHyOz

equation : CxHyOz + O2 = CO2 +H2O

(5.015). (7.35) (2.99)

mass of C =12g/mol * 7.35/44g/mol= 2.005g

mass of H =2g/mol* 2.99g/18g/mol =0.333

mass of O2 in CxHyOz = 5.015-(2.005+0.333)= 2.678

moles of Carbon = 2.005/12= 0.167

moles of Hydrogen = 0.333/1 = 0.333

moles of oxygen = 2.678/16 = 0.168

divide all by smallest moles value = C(0.500) H(2) O(0.500)

multiply by whole number (2) = C(1) H(2) O(1)

empirical formula =CH2O.